OpenStudy (anonymous):

Simplify. b3·b-5·b11 A) b9 B) b15 C) b19 D) b-165

OpenStudy (diyadiya):

\[b^3.b^{-5}.b^{11}\]?

OpenStudy (hoblos):

you mean (b^3)(b^-5)(b^11) ?

OpenStudy (hoblos):

if so .... add the powers

OpenStudy (diyadiya):

\[\huge x^m \times x^n = x^{m+n}\] Use this !

OpenStudy (anonymous):

soo what is it

OpenStudy (diyadiya):

What did you get ?

OpenStudy (anonymous):

its A. b^9

OpenStudy (diyadiya):

\[\large b^{3+(-5)+11}= answer \]